# Criticality-Guided Efficient Pruning in Spiking Neural Networks Inspired by Critical Brain Hypothesis

### v2

@@ Abstract
Spiking Neural Networks (SNNs) have gained considerable attention due to the energy-efficient and multiplication-free characteristics. The continuous growth in scale of deep SNNs poses challenges for model deployment. Network pruning reduces hardware resource requirements of model deployment by compressing the network scale. However, existing SNN pruning methods cause high pruning costs and performance loss because the pruning iterations amplify the training difficulty of SNNs. In this paper, inspired by the critical brain hypothesis in neuroscience, we propose a regeneration mechanism based on the neuron criticality for SNN pruning to enhance feature extraction and accelerate the pruning process. Firstly, we propose a low-cost metric for the criticality in SNNs. Then, we re-rank the pruned structures after pruning and regenerate those with higher criticality to obtain the critical network. Our method achieves higher performance than the current state-of-the-art (SOTA) method with up to 95.26% reduction of pruning cost. Moreover, we investigate the underlying mechanism of our method and find that it efficiently selects potential structures and learns the consistent feature representation.
@@ 1 Introduction
Spiking Neural Networks (SNNs) have received considerable attention in recent years (Wu et al., 2019; Zheng et al., 2021) as the third generation of neural networks (Maass, 1997). SNNs are widely employed in resource-constrained hardware (Akopyan et al., 2015; Davies et al., 2018; Pei et al., 2019), relying on energy-efficient and multiplication-free advantages from sparse spike signals.
@@
The limited computational and storage resources of hareware pose challenges to implementing deep SNNs with large-scale parameters. Network pruning provides a potential solution to address resource issues. There have been a lot of efforts to unstructured pruning for SNNs. Methods inspired by the human brain include modeling synaptic regeneration processes (Kundu et al., 2021) and dendritic motion (Kappel et al., 2015). Qi et al. (2018) designed connection gates during the training process for synaptic pruning. Recent efforts reference techniques from Artificial Neural Network (ANN) pruning. Kim et al. (2022) explored the ”lottery ticket” hypothesis in SNN pruning with Iterative Magnitude Pruning (IMP). Deng et al. (2021) combined Spatio-Temporal Backpropagation (STBP) and Alternating Direction Method of Multipliers (ADMM) and proposed an activity-based regularization method. Grad R (Chen et al., 2021) improved the Deep R method by introducing a weight regeneration mechanism. RCMO-SNN (Chen et al., 2023) proposed an end-to-end Minimax optimization method for sparse learning.
@@
However, current state-of-the-art (SOTA) pruning methods for SNNs cause high pruning costs and performance loss. Table 1 summarizes the required epochs and pruned accuracy on ResNet19 with about 95% sparsity for CIFAR100 using different pruning methods. Existing SOTA methods require over 2000 epochs (Chen et al., 2021; Kim et al., 2022; Chen et al., 2023). Huang et al. (2023) develop NDSNN based on sparse training, reducing pruning cost to 300 epochs but with a significant loss of accuracy. Other efforts to reduce pruning costs include compressing timesteps during the pruning process (Kim et al., 2022, 2023). However, these efforts only achieve acceleration ratios of 1.12 to 1.59 and lead to significant accuracy losses.
@@
The high cost of pruning comes from the expensive pruning iterations (Chen et al., 2021; Kim et al., 2022; Chen et al., 2023). Multiple rounds of pruning and fine-tuning are employed in SOTA SNN pruning methods to achieve high-performance pruned models. However, the utilization of sparse spike signals in SNNs introduces problems like spike and gradient vanishing, which hinder both forward and backward propagation during SNN training, consequently leading to inefficient information passing and feature extraction (Zheng et al., 2021; Wu et al., 2019; Shrestha et al., 2021). The iterative pruning exacerbates this problem, requiring the long recovery time after each pruning round (Kim et al., 2022; Chen et al., 2023), or forcing magnitude reduction of single pruning step for rapid recovery (Chen et al., 2021). Both lead to a multiplication of pruning costs, still accompanying a loss in accuracy.
In this paper, we are the first to leverage criticality to improve feature extraction and achieve high-efficiency pruning for SNNs. The critical brain hypothesis (Turing, 2009) reveals that brain neurons with the critical state are highly sensitive to inputs and exhibit enhanced information transmission, associated with the complex feature extraction and learning process of the brain (Kinouchi and Copelli, 2006; Beggs, 2008; Shew et al., 2009; Beggs and Timme, 2012). Inspired by this, our goal is to leverage the advantages of criticality to enhance feature extraction and speed up the pruning process in SNN pruning. Based on the association between criticality and the barely-excitable state of a single neuron (Gal and Marom, 2013), we propose a metric for neuron criticality and further relate it to the surrogate function derivative to avoid additional computational costs. We then design a regeneration mechanism based on neuron criticality to introduce criticality into the pruning and fine-tuning iterations and preserve critical neurons. The proposed method is evaluated on unstructured and structured pruning and achieves SOTA performances. Moreover, we explore the underlying impact of our method and find that it efficiently selects potential structures and improves the uniformity of features.
@@
  • Inspired by the critical brain hypothesis in neuroscience, we propose a metric for the neuron criticality of SNNs and design a regeneration mechanism to improve feature extraction and achieve efficient pruning for SNNs.
@@
  • We explore the underlying impact of our method and find that: (1) our method selects critical structures with latent potential, which become more important after fine-tuning. (2) Our method improves the uniformity of class features between the train and test samples and reduces noise. These improve the stability and reduce the overfitting during the recovery phase.
@@ 2.1.1 Pruning Criterion
2.1.2 Pruning Schedule
The setting of pruning schedules affects the performance of pruned networks. Iterative pruning follows the pruning-fine-tuning pattern with settings such as the pruning ratio in one iteration, pruning interval, and number of iterations. Zhu and Gupta (2017) proposed a gradual sparse ratio schedule with a unified pruning interval for unstructured pruning and was followed by subsequent studies (Gale et al., 2019; Liu et al., 2021). Molchanov et al. (2016, 2019); Ding et al. (2021) implemented the iterative pruning on channel level. Frankle and Carbin (2018) proposed the lottery hypothesis with a fixed pruning ratio and a longer pruning interval while inheriting the surviving parameters of the initial model after each iteration. Kim et al. (2022) proved the lottery hypothesis in SNN pruning. One-shot pruning refers to achieving the target sparsity once. You et al. (2019) proposed using the Hamming distance of pruning masks in the early stages of training to determine the pruning time. Kim et al. (2022) implemented the above method on SNN pruning.
(a) Unstructured Pruning
(b) Structured Pruning
@@ 3.1 Spiking Neural Network
In SNN, a neuron receives signals from model inputs or pre-synaptic neurons and modulates the membrane potential. When the membrane potential exceeds the threshold, An output spike is generated and sent to the post-synaptic neuron. The fire process is described as:
@@ 4.1 Criticality from Biology to SNNs
Inspired by the critical brain hypothesis , We aim to exploit the advantages of the critical state to improve SNN pruning. Specifically, we hope that pruned models retain more critical characteristics and maintain sensitivity to inputs and efficiency in feature extraction, ultimately accelerating pruning and improving post-pruning performance. We start from a single neuron to exploit the criticality in SNNs since it shows the lowest level of criticality in neuroscience (Gal and Marom, 2013). Previous works (Hu et al., 2016; Liu et al., 2022) have improved pruning through properties exhibited by neurons during training and pruning processes. Drawing from these works, a natural idea is to compute the criticality score of neurons and select ones with higher criticality.
@@
The key challenge is to find a suitable metric for neuron criticality. Gal and Marom (2013) reveal that criticality is correlated with the barely-excitable state of neurons. Plesser and Gerstner (2000); Maass (2014) indicate that the magnitude of the membrane potential is positively correlated with the excitation probability. As the membrane potential approaches the threshold voltage, the excitation probability rapidly transitions between 0 and 1. Following this line, we propose that neuronal criticality is related to the distance between the membrane potential and the threshold voltage. The behavior of a neuron is considered to have higher criticality when the membrane potential is closer to the threshold voltage. Considering pruning and training processes, we suggest that the derivative of the surrogate function, such as $g^{\prime}$ in Eq.2b, serves as a criticality metric. It offers several advantages: (1) The derivatives of mainstream surrogate functions reach the highest value when the membrane potential equals the threshold voltage and rapidly decreases with increasing distance, effectively reflecting the criticality changes of a neuron’s behavior. (2) The derivative of the surrogate function can be directly obtained during model training with minimal additional computational cost. (3) The derivative of the surrogate function has a unified value range, enabling global ranking.
4.2 Regeneration via Neuron Criticality
We aim to develop a mechanism compatible with various pruning strategies to preserve critical neurons. To achieve this, we combine the criticality metric mentioned above with a neuron regeneration mechanism. The regeneration mechanism is inspired by synaptic regeneration in the brain and used to reactivate pruned structures. Gradient magnitude has been used as a criterion for regeneration in previous ANN pruning methods like RigL (Evci et al., 2020) and GraNet (Liu et al., 2021), but it lacks motivation and biological interpretability for SNNs. We implement a regeneration mechanism based on neuron criticality to maintain the critical state of the pruned model.
@@ 4.2 Regeneration via Neuron Criticality
Specifically, to avoid the sparsity of the model after regeneration from being lower than the target, we first temporarily expand the target sparsity at the current pruning iteration $s_{t}$ to $s_{t}^{\prime}$:
@@
where $S$ is the set of surviving pruning structures after the pruning iteration. $S’$ ’ is the set of all pruned structures. $C(x)$ denotes the criticality score of $x$. To compute the criticality score, we have
@@
Eq.5a and 5b illustrate the approach to obtaining criticality scores for a neuron $e$. $u_{e,i}$ represents the membrane potential of $e$ at $i$th time step. $T$ is the total number of time steps. $g^{\prime}$ is the derivative of the surrogate function. $N$ is the total number of samples used to compute the criticality score. $aggregate$ represents the way to aggregate the results for the corresponding pruning structure. For the fully connected layer, we directly compute the average criticality score for each neuron. For convolutional layers, we explore two approaches: mean aggregation and max aggregation. We finally employ max operation due to its superior performance in the experiments. We suppose that max aggregation preserves the highest response within the corresponding region, preventing robust responses from being diminished by neighboring elements within the feature map.
4.3 Overall Pruning Strategy
So far, we have known how to calculate the criticality scores of neurons. In order to evaluate the impact of preserving critical neurons on sparse SNN, we perform pruning on SNNs at both unstructured (connection) and structured (channel) levels. As depicted in Figure 1, our approach employs the basic pruning techniques and incorporates the regeneration mechanism based on neuron criticality. We introduce the workflow below. Implementation details are contained in Appendix A.

### v3

# Brain-Inspired Efficient Pruning: Exploiting Criticality in Spiking Neural Networks

@@ Abstract
Spiking Neural Networks (SNNs) have gained significant attention due to the energy-efficient and multiplication-free characteristics. Despite these advantages, deploying large-scale SNNs on edge hardware is challenging due to limited resource availability. Network pruning offers a viable approach to compress the network scale and reduce hardware resource requirements for model deployment. However, existing SNN pruning methods cause high pruning costs and performance loss because they lack efficiency in processing the sparse spike representation of SNNs. In this paper, inspired by the critical brain hypothesis in neuroscience and the high biological plausibility of SNNs, we explore and leverage criticality to facilitate efficient pruning in deep SNNs. We firstly explain criticality in SNNs from the perspective of maximizing feature information entropy. Second, We propose a low-cost metric for assess neuron criticality in feature transmission and design a pruning-regeneration method that incorporates this criticality into the pruning process. Experimental results demonstrate that our method achieves higher performance than the current state-of-the-art (SOTA) method with up to 95.26% reduction of pruning cost. The criticality-based regeneration process efficiently selects potential structures and facilitates consistent feature representation.
keywords:
Spiking neural network , Brain-inspired computing , Network pruning
journal: Neural Networks
[…]
@@ 1 Introduction
Spiking Neural Networks (SNNs) have received significant attention in recent years (Wu et al., 2019; Zheng et al., 2021) as the third generation of neural networks (Maass, 1997). SNNs are widely employed in resource-constrained hardware (Akopyan et al., 2015; Davies et al., 2018; Pei et al., 2019; Liu et al., 2023), relying on energy-efficient and multiplication-free advantages from sparse spike signals.
The limited computational and storage resources of hareware pose challenges to implementing deep SNNs with large-scale parameters. Network pruning provides a potential solution to address resource issues. There have been a lot of efforts to unstructured pruning for SNNs. Methods inspired by the human brain include modeling synaptic regeneration processes (Kundu et al., 2021) and dendritic motion (Kappel et al., 2015). Recent efforts reference to techniques from Artificial Neural Network (ANN) pruning. Kim et al. (2022) explored the “lottery ticket” hypothesis in SNN pruning with Iterative Magnitude Pruning (IMP). Deng et al. (2021) combined Spatio-Temporal Backpropagation (STBP) and Alternating Direction Method of Multipliers (ADMM) and proposed an activity-based regularization method. Grad R (Chen et al., 2021) improved the Deep R method by introducing a weight regeneration mechanism. RCMO-SNN (Chen et al., 2023) proposed an end-to-end Minimax optimization method for sparse learning.
@@
However, current state-of-the-art (SOTA) pruning methods spend significant time costs in the SNN pruning process compared to ANNs. Table 1 summarizes the required epochs and pruned accuracy on ResNet19 for CIFAR100 using different pruning methods. Existing SOTA methods  (Chen et al., 2021; Kim et al., 2022; Chen et al., 2023) require over 2000 epochs to achieve approximately 95% sparsity on ResNet19. Efforts to reduce pruning costs include NDSNNs based on sparse training (Huang et al., 2023), which reduce the epoch requirement to 300 but with a significant loss of accuracy. Other approaches  (Kim et al., 2022, 2023) attempt to reduce pruning costs by compressing time steps during the pruning process, but these methods only achieve limited acceleration ratios (1.12 to 1.59) and still result in significant performance losses.
@@
The high pruning cost comes from the multiple rounds of pruning and long fine-tuning time required by SOTA SNN pruning methods to achieve high-performance pruned models. Unlike the continuous value processing in ANNs, SNNs transmit information using sparse spike signals, relying on specialized neuron activations and neural signal encoding, which poses challenges for learning spike feature patterns in deep SNNs (Zheng et al., 2021; Wu et al., 2019; Shrestha et al., 2021). The multiple rounds of pruning and fine-tuning exacerbate the learning challenge. Specifically, it requires the long recovery time after each pruning round (Kim et al., 2022; Chen et al., 2023), or forces a reduction in the magnitude of single pruning step for rapid recovery (Chen et al., 2021). Both lead to a multiplication of pruning costs, accompanying a loss in accuracy. Therefore, it is necessary to develop a new SNN pruning method that can effectively preserve the critical spike patterns and reduce fine-tuning costs without sacrificing performance.
In this paper, inspired by the critical brain hypothesis in neuroscience, we explore and leverage criticality to facilitate efficient pruning in deep SNNs. The critical brain hypothesis (Turing, 2009) suggests that neurons operating at a critical state possess excellent information transmission capabilities, crucial for complex feature extraction and learning in the brain (Kinouchi and Copelli, 2006; Beggs, 2008; Shew et al., 2009; Beggs and Timme, 2012). SNNs mimic the mechanisms of biological neural systems and demonstrate high biological plausibility. So we employ the criticality in SNN pruning, aiming to maximally preserve the feature extraction capability in pruned models and accelerate the entire pruning process.
We first analyze criticality in SNNs through information entropy, finding that neurons at critical states can encode and transmit information more efficiently. By combining the principle of maximizing feature information entropy and observations on the behavior of neurons at critical states (Gal and Marom, 2013), we propose a criticality metric based on the derivative of neuron surrogate functions. This metric assesses the criticality of neurons in information processing while avoiding additional computational costs. To integrate this criticality into the pruning process of SNNs, we design a pruning-regeneration method. Experimental results show that our method, evaluated on both unstructured and structured pruning, demonstrates higher performance and lower pruning costs compared with state-of-the-art methods. The criticality metric effectively identifies and preserves critical structures. Our method achieves lower intra-cluster variance within the class and higher feature cosine similarity between the training samples and test samples, confirming the effectiveness of the criticality mechanism in enhancing feature extraction and consistency of pruned models.
@@
  • 1.
    Inspired by the critical brain hypothesis in neuroscience, we propose a metric for the neuron criticality of SNNs and design a pruning-regeneration method to improve feature extraction and achieve efficient pruning for SNNs.
  • 2.
    We verify the utility of neuron criticality and its metric in SNNs from the perspective of maximizing feature information entropy. Furthermore, Experimental results show that the proposed criticality metric efficiently selects potential structures, reduces intra-cluster variance in 93% of classes on average, and increases feature cosine similarity between training and test samples in over 98% of classes. These findings highlight improved feature extraction and consistency in pruned models.
  • 3.
@@ 2.1.1 Pruning Criterion
2.1.2 Pruning Strategy
Pruning strategy affects the performance of pruned networks. Iterative pruning follows the pruning-fine-tuning pattern with settings such as the pruning ratio in one iteration, pruning interval, and number of iterations. Zhu and Gupta (2017) proposed a gradual sparse ratio schedule with a unified pruning interval for unstructured pruning and was followed by subsequent studies (Gale et al., 2019; Liu et al., 2021). Molchanov et al. (2016, 2019); Ding et al. (2021) implemented the iterative pruning on channel level. Frankle and Carbin (2018) proposed the lottery hypothesis with a fixed pruning ratio and a longer pruning interval while inheriting the surviving parameters of the initial model after each iteration. Kim et al. (2022) proved the lottery hypothesis in SNN pruning. One-shot pruning refers to achieving the target sparsity once. You et al. (2019) proposed using the Hamming distance of pruning masks in the early stages of training to determine the pruning time. Kim et al. (2022) implemented the above method on SNN pruning.
@@ 3.1 Spiking Neural Network
In SNNs, a neuron receives signals from model inputs or pre-synaptic neurons and modulates the membrane potential. When the membrane potential exceeds the threshold, An output spike is generated and sent to the post-synaptic neuron. The fire process is described as:
@@ 4.1 Criticality from Biology to SNNs
We aim to exploit criticality in SNN pruning to improve feature extraction capability in pruned models and accelerate the entire pruning process. To analyze criticality in SNNs, we employ information entropy and correlate criticality with the maximization of feature information entropy.
4.1.1 Criticality of Single Neuron
We start from a single neuron to analyze the criticality in SNNs since it is the the fundamental unit exhibiting criticality in neuroscience (Gal and Marom, 2013). Gal and Marom (2013) reveal that criticality is correlated with the barely-excitable state of neurons. Plesser and Gerstner (2000); Maass (2014) indicate that the magnitude of the membrane potential is positively correlated with the excitation probability. As the membrane potential crosses the threshold voltage, the firing probability rapidly increases from 0 to 1. Following this line, we propose that neuronal criticality is related to the distance between the membrane potential and the threshold voltage. The behavior of a neuron is considered to have higher criticality when the membrane potential is closer to the threshold voltage.
To further explore the relationship between criticality and neuronal dynamics, information entropy is employed to quantify the information transmission capacity of neurons and SNNs. We compute and maximize the information entropy of the neuron’s output. The information entropy of the output of a single neuron in one time step is defined as
$H(X)=-\sum_{x\in\mathcal{X}}{p(X=x)\ln{p(X=x)}},$ (3)
where the random variable $X$ represents the neuron’s output, with a value set of $\{0,1\}$. We denote the firing probability $p(X=1)$ as $p_{f}$, and Eq. 3 can be reformulated as
$H(X)=-p_{f}\ln{p_{f}}-(1-p_{f})\ln{(1-p_{f})}.$ (4)
Eq. 4 is a concave function with respect to $p_{f}$, achieving its maximum value of $\ln{2}$ when $p_{f}=0.5$.
We introduce $p(u)$ to represent the neuron firing probability $p_{f}$, where $u$ is the membrane potential of the neuron at the current time step, decoupling the firing probabilities across different time steps. As previously mentioned, experiments have demonstrated a rapid increase in firing probability from 0 to 1 as the membrane potential increases. And $p(u)=0.5$ is achieved when the membrane potential is equal to the threshold voltage. Since Eq. 4 is concave, the information entropy of the neuron’s output is greater when the membrane potential is closer to the threshold voltage.
4.1.2 Criticality of Neuron Population
For populations of neurons, we define the joint probability distribution of the output feature map of a layer as $p(\mathbf{X})$. The information entropy of the feature map output can be represented as
$H(\mathbf{X})=-\sum_{\mathbf{x}\in\mathbf{\mathcal{X}}}p(\mathbf{X}=\mathbf{x})\ln{p(\mathbf{X}=\mathbf{x})},$ (5)
where $\mathbf{X}$ is a random vector representing the values of feature maps. As the size of the feature maps increases, the sample space of $\mathbf{X}$ grows exponentially, making it difficult to directly analyze the maximum value of $H(\mathbf{X})$. In fact, we estimate the joint entropy $H(\mathbf{X})$ using the sum of the individual neuron entropies $\sum_{i=1}^{n}{H(X_{i})}$. We have
$\sum_{i=1}^{n}H(X_{i})-H(\mathbf{X})$ (6)
$=\sum_{\mathbf{x}\in\mathbf{\mathcal{X}}}p(\mathbf{x})\ln p(\mathbf{x})-\sum_{i=1}^{n}\sum_{a\in\mathcal{X}_{i}}p(X_{i}=a)\ln p(X_{i}=a)$ (7)
$=\sum_{\mathbf{x}\in\mathbf{\mathcal{X}}}p(\mathbf{x})\ln p(\mathbf{x})-\sum_{i=1}^{n}\sum_{a\in\mathcal{X}_{i}}\sum_{\mathbf{x}\in\mathbf{\mathcal{X}}}p(\mathbf{x})I_{x_{i}=a}(\mathbf{x})\ln p(X_{i}=a)$ (8)
$=\sum_{\mathbf{x}\in\mathbf{\mathcal{X}}}p(\mathbf{x})\ln p(\mathbf{x})-\sum_{\mathbf{x}\in\mathbf{\mathcal{X}}}p(\mathbf{x})\ln{\prod_{i=1}^{n}p(X_{i}=x_{i})}$ (9)
$=\sum_{\mathbf{x}\in\mathbf{\mathcal{X}}}p(\mathbf{x})\ln{\frac{p(\mathbf{x})}{\prod_{i=1}^{n}p(X_{i}=x_{i})}}.$ (10)
In Eq. 8, we expand $p(X_{i}=a)$ with $\mathbf{x}\in\mathbf{\mathcal{X}}$. $I_{x_{i}=a}(\mathbf{x})$ is an indicator function, which is 1 only when $x_{i}=a$ and 0 otherwise. The difference between the full entropy and the sum of individual neuron entropies is transformed into Eq. 10. The inequality follows from the fact that Eq. 10 represents the Kullback-Leibler (KL) divergence between the probability distributions $p(\mathbf{x})$ and $\prod_{i=1}^{n}p(X_{i}=x_{i})$.The KL divergence indicates the distance between $H(\mathbf{X})$ and $\sum_{i=1}^{n}{H(X_{i})}$ and the distance is 0 when $p(\mathbf{x})=\prod_{i=1}^{n}p(X_{i}=x_{i})$.
We similarly introduce $p(u_{i})$ to represent $p(X_{i}=1)$, where $u_{i}$ is the membrane potential of the $i$-th neuron at the current time step. The distribution of membrane potentials is reorganized through the Batch Normalization (BN) during the forward propagation process, reducing the correlation of firing rate $p(u_{i})$ between neurons and making $p(\mathbf{x})$ close to $\prod_{i=1}^{n}p(X_{i}=x_{i})$. On this basis, we estimate $H(\mathbf{X})$ using $\sum_{i=1}^{n}H(X_{i})$, which has been proven to be reasonable during the maximization process (Dayan and Abbott, 2005). From Eq. 4, it is easy to obtain that $\sum_{i=1}^{n}H(X_{i})$ is concave in $R^{n}$, and it takes the maximum value when $p(u_{i})=0.5,\forall i$. As mentioned in Section 4.1.1, $p(u_{i})$ rapidly increases from 0 to 1 when $u_{i}$ crosses the threshold voltage. $p(u_{i})$ reaches $0.5$ at a point where $u_{i}$ is infinitely close to the threshold voltage. In this case, the neuron approaches a critical state, transmitting a greater amount of information at both the single neuron and feature map levels.
4.2 Metric of Neuron Criticality
So far, we have known that when the membrane potential is closer to the threshold voltage, the neuron’s criticality is higher and the output feature carries more information. To integrate criticality into SNN pruning, it is necessary to identify an appropriate metric to assess the criticality of neurons. The primary challenge lies in capturing the rapid changes of criticality near the threshold voltage and ensuring consistent, low-cost computation during model training.
@@ 4.2 Metric of Neuron Criticality
We propose that the derivative of the surrogate function, such as $g^{\prime}$ in Eq. 2b, can serve as a criticality metric. It offers several advantages: (1) The derivatives of widely-used surrogate functions reach the peak when the membrane potential equals the threshold voltage and rapidly decreases with increasing distance, effectively reflecting the criticality changes of a neuron’s behavior. (2) The derivative of the surrogate function can be directly obtained during model training without additional computational cost. (3) The derivative of the surrogate function has a unified value range, enabling global ranking.
(a) Unstructured Pruning
(b) Structured Pruning
4.3 Overall Pruning Strategy
4.3.1 Regeneration via Criticality
Our goal is to develop a mechanism compatible with exist pruning strategies to leverage both the criticality and the established advantages of SOTA methods. To achieve this, we combine the criticality metric mentioned above with a neuron regeneration mechanism. The regeneration mechanism, inspired by neuroregeneration in the nervous system where new neurons and connections are synthesized to recover damage, allows the pruned network to regenerate lost connections, thereby improving performance after pruning.
@@ 4.3.1 Regeneration via Criticality
We implement a regeneration mechanism based on neuron criticality to maintain the critical state of the pruned model. Specifically, to avoid the sparsity of the model after regeneration from being lower than the target, we first temporarily expand the target sparsity at the current pruning iteration $s_{t}$ to $s_{t}^{\prime}$:
@@
where $S$ is the set of surviving pruning structures after the pruning iteration. $S’$ ’ is the set of all pruned structures in current iteration. $C(x)$ denotes the criticality score of $x$. To compute the criticality score, we have
@@
Eq. 13a and 13b illustrate the approach to obtaining criticality scores for a neuron $e$. $u_{e,i}$ represents the membrane potential of $e$ at $i$th time step. $T$ is the total number of time steps. $g^{\prime}$ is the derivative of the surrogate function. $N$ is the total number of samples used to compute the criticality score. $aggregate$ represents the way to aggregate the results for the corresponding pruning structure.
For the fully connected layer, we directly compute the average criticality score for each neuron. For convolutional layers, we explore two approaches: mean aggregation and max aggregation. We finally employ max operation due to its superior performance in the experiments. We consider that max aggregation preserves the highest response within the corresponding region, preventing robust responses from being diminished by neighboring elements within the feature map.